On January 2, 2019, Gino Co. issued at par $200,000 of 9% convertible bonds. Each $1,000 bond is convertible into 20 shares. No bonds were converted during 2019. Gino had 50,000 shares of common stock outstanding during 2019. Gino's 2019 net income was $150,000 and the income tax rate was 25%. Compute Gino's EPS and diluted earnings per share for 2019 (rounded to the nearest penny). Show all computations
Answer #1
a.) Earnings for Basic EPS equals to Net Income        150,000
Shares for Basic EPS equals to Common stock outstanding            50,000
Basic EPS $ 3.00 =150000/50000
(Earnings / Shares )
b.) Net Income        150,000
Add: Interest on Bond, Net of tax            13,500 =200000*9%*(1-25%)
Earnings for Diluted EPS $ 163,500
Number of Common stock            50,000
Add: Number of Common stock convertible from Bond              4,000 =200000*20/1000
Shares for Diluted EPS           54,000
Diluted earning per share for 2019 $ 3.03 =163500/54000
(Earnings / Shares )